OpenStudy (anonymous):

A projectile is fired in such a way that its horizontal range is equal to 13.5 times its maximum height. What is the angle of projection?

OpenStudy (anonymous):

R tan (theta) =4H

OpenStudy (anonymous):

\[R =( u^2 \sin2\theta)/g \] \[H = (u^2 \sin^2\theta)/2g\] Put R = 13.5 H You'll get, \[\tan \theta = 4/13.5\] \[\theta = 16.504\] Is it correct ?

OpenStudy (anonymous):

(R/H) =13.5=4 cotθ => θ =16.504

OpenStudy (anonymous):

yes it was right!! thank you so much!!!

OpenStudy (anonymous):

r=13.5h =>u^2=20h/sin^2a =>r=u^2sin a.cos a/10 =>a=cot^(-1)(13.5/4)

OpenStudy (anonymous):

:)